OpenStudy (anonymous):

ALG 2 A football is kicked into the air from an initial height of 3 feet. The height, in feet, of the football above the ground is given by s(t) - -16t^2+35t + 3 , where t is time in seconds and>0 . Which is closest to the time when the football will be 20 feet above the ground?

OpenStudy (anonymous):

2.72 seconds 0.73 seconds or 1.46 seconds 0.53 seconds 0.53 seconds or 2.72 seconds

OpenStudy (michele_laino):

here we have to solve this equation: \[\huge - 16{t^2} + 35t + 3 = 20\]

OpenStudy (michele_laino):

after a simplification, we get: \[\huge 16{t^2} - 35t + 17 = 0\] next, please apply the standard formula: \[\huge t = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]